OpenStudy (anonymous):

HELP MEDAL + FAN 22 : 66 What is the above ratio in lowest terms

OpenStudy (anonymous):

@angel12310

OpenStudy (jadedry):

Divide both 22 and 66 by 22

OpenStudy (kweenlexis):

22 / 66 = (2 * 11) / (2 * 3 * 11) = ((2 * 11) : (2 * 11)) / ((2 * 3 * 11) : (2 * 11)) = (22 : 22) / (66 : 22) = 1 / 3

OpenStudy (anonymous):

yes its 1:3

OpenStudy (kweenlexis):

/ suppose 2 be :

OpenStudy (anonymous):

Thanks

OpenStudy (kweenlexis):

you welcome!